I have a problem on my homework that is .02(300)-x=15-.05x and the answer is supposed to be 180 but i keep getting -9. i don't understand how you get 180. my friends get -9 also. thank you for your help :)
Answer by Alan3354(69443)

I have a problem on my homework that is .02(300)-x=15-.05x and the answer is supposed to be 180 but i keep getting -9.
-----
.02(300)-x=15-.05x
Multiply by 100 to eliminate fractions
2*300 - 100x = 1500 - 5x
600 - 100x = 1500 - 5x
Add 100x
600 = 1500 + 95x
Subtract 1500
-900 = 95x
x = -95/900 = -19/180
--------
How did you get -9 ?